Find the equation of a parabola that has a vertex of (9,13) and a y-intercept of 94. Write answer in general form.
Answer by lwsshak3(6509)

Find the equation of a parabola that has a vertex of (9,13) and a y-intercept of 94. Write answer in general form.
**
Standard form of equation for parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
..
Plug in given coordinates of vertex and point (0,94), and solve for A.
94=A(0-9)^2+13
94=81A+13
81=81A
A=1
Equation:
y=(x-9)^2+13